$SPAD/src/input drawcurv.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all
-- a circle
seg1 : SEG FRAC INT := -3/2..3/2
range1 : LIST SEG FRAC INT := [seg1,seg1]
draw(x**2 + y**2 = 1,x,y,range == range1,title == "Unit Circle")
-- ditto
seg2 : SEG FLOAT := -1.1..1.1
range2 : LIST SEG FLOAT := [seg2,seg2]
draw(x**2 + y**2 = 1,x,y,range == range2,title == "Unit Circle, 2nd Graph")
-- an ellipse
seg3 : SEG FRAC INT := -4..4
range3 : LIST SEG FRAC INT := [seg3,seg3]
draw(x**2/9 + y**2/4 = 1,x,y,range == range3,_
                             toScale == true,title == "Ellipse")
-- another ellipse
seg4 : SEG FRAC INT := -5..5
range4 : LIST SEG FRAC INT := [seg4,seg4]
draw(5*x**2 - 8*x*y + 5*y**2 = 9,x,y,range == range4)
-- a parabola
seg5a : SEG FRAC INT := -3..14
seg5b : SEG FRAC INT := -3..10
range5 : LIST SEG FRAC INT := [seg5a,seg5b]
draw(x**2 - 2*x*y + y**2 - x - 2 = 0,x,y,range == range5)
-- a hyperbola

draw(x**2/4 - y**2 = 1,x,y,range == range4)
-- an elliptic curve
seg6a : SEG FRAC INT := -2..2
seg6b : SEG FRAC INT := -2..1
range6 : LIST SEG FRAC INT := [seg6a,seg6b]
draw(y**2 + y = x**3 - x,x,y,range == range6)
-- ovals
seg7a : SEG FRAC INT := -8..8
seg7b : SEG FRAC INT := -4..4
range7 : LIST SEG FRAC INT := [seg7a,seg7b]
eq1 := (x**2 + y**2 + 5**2)**2 - (6**4 + 4*5**2*x**2) = 0
draw(eq1,x,y,range == range7,title == "Cassinian oval with one loop")
seg8a : SEG FRAC INT := -10..10
range8 : LIST SEG FRAC INT := [seg8a,seg7b]
eq2 := (x**2 + y**2 + 7**2)**2 - (6**4 + 4*7**2*x**2) = 0
draw(eq2,x,y,range == range8,title == "Cassinian oval with two loops")

———-

References^{[1] nothing}

5